OpenStudy (anonymous):

A Ferris wheel car moves from point C to point D on the circle shown below:

OpenStudy (anonymous):

What is the arc length the car traveled, to the nearest hundredth? 8.30 feet 11.30 feet 12.41 feet 13.15 feet Can you please help me with this one? @zaibali.qasmi

OpenStudy (anonymous):

@zaibali.qasmi

OpenStudy (anonymous):

I need some help on this one too if you can help :) The volume of a pyramid that fits exactly inside a cube is 49 cubic feet. What is the volume of the cube? 7 cubic feet 147 cubic feet 241 cubic feet 2401 cubic feet

OpenStudy (anonymous):

what about the first question?

OpenStudy (anonymous):

@zaibali.qasmi

OpenStudy (anonymous):

are you able to help with that question so i can move on?

OpenStudy (anonymous):

use this...

OpenStudy (anonymous):

thank you so much one last one and i promise i'm done asking for help lol

OpenStudy (anonymous):

Using the following equation, find the center and radius: x2 + 4x + y2 − 6y = −4 The center is located at (−2, 3), and the radius is 3. The center is located at (2, −3), and the radius is 3. The center is located at (−2, 3), and the radius is 9. The center is located at (2, −3), and the radius is 9.

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

anyone's help would be great!

OpenStudy (anonymous):

thank you so so so much for all your wonderful help. And i truly apologize for bothering

OpenStudy (anonymous):

see how just complte the squares for x^2 and y^2 separately as here below.. (x^2+2(x)(2)+(2)^2)-2^2+(y^2-2y(3)+3^2)-3^2=-4 =>(x+2)^2+(y-3)^2-4-9=-4 =>(x+2)^2+(y-3)^2=9=3^2 so here is (-2,3) and centre is 3

OpenStudy (anonymous):

aahhhh I see what you did thank you very much!